Two electric charges of 9uC and -3uC are placed 0.16 m apart in air . there be a point at which electric potential is zero on the line joining the two charges and in between them . The distance of P from 9uC charge is? A) 0.14 m B) 0.12 m C) 0.08 m D) 0.06 m
1
Expert's answer
2013-04-22T10:39:55-0400
To find this point we must write condition of equality of potential from two charges. Let 9uC be located at 0 on x axis
then, this condition can be written as

kq1/p^2 = kq2/(0.16-p)^2

where q1 and q2 are charges, and p is distance from origin to equal potentials-point

so we find

9/p^2 = -3/(0.16^2-0.32p+p^2)

3p^2-9*0.08p+9*0.0064 = 0

p=0.12

correct answer is
B) 0.12 m